GTS meeting 6 20 16
- 1 mA to the dump! at 275 kV
- The QE was estimated to be around 0.3%
- We discovered that the Haimson coil corrector was not steering the beam properly. Later on Joe Grames and Yan Wang discovered it was wired up as a quad. They corrected the problem.
- Beam on Viewers
- Adriana and Shukui did a calibration of the laser power meter located inside the hutch and read in the control room with respect to the laser power going into the gun vacuum chamber.
- Did the cathode got damaged here?
- Image of cathode damage
- Laser and electron beam after discovering bad laser shape
- Adriana discovered that the bad laser shape is due to a damaged polarizer in the hutch. Shukui gave permission to fix the problem.
- The graph below shows the QE curves for various runs, indicating QE drop with running time.
- Mamun and Chio have made a 12% multi-alkali in Nb substrate